Could someone explain this to me? thanks so much!
A building casts a 52-foot shadow at the same time taylor casts a 4-foot shadow. Taylor and her shadow form a triangle that is similiar to the triangle formed by the building and it's shadow. If Taylor is 5 feet tall, what is the height h of the building?

Answer by venugopalramana(3286)

A building casts a 52-foot shadow at the same time taylor casts a 4-foot shadow. Taylor and her shadow form a triangle that is similiar to the triangle formed by the building and it's shadow. If Taylor is 5 feet tall, what is the height h of the building?
SINCE IT IS GIVEN THAT TAYLOR AND THE BULDING AND THEIR SHADOWS FORM SIMILAR TRIANGLES...THEIR SIDES WILL BE IN PROPORTION.SO
SHADOW OF TAYLOR/HER HEIGHT=SHADOW OF BUILDING /ITS HT.
4/5=52/H
H=52*5/4=65'